# Heterogeneous Treatment Effects and

### v3

@@ 6.2 Improving the interpretation of HTEs as mechanism tests
This framework provides guidance for the interpretation of HTEs when researchers are trying to make inferences about which causal mechanisms are active. First, while the presence of HTEs provide evidence that a mechanism is active when the exclusion assumption holds and an outcome is directly affected, the absence of HTEs is less informative (even under the same exclusion assumptions). In this sense, a lack of HTEs cannot be used to “rule out” a candidate mechanism or show that it is inert. As Proposition 2 shows, even when an outcome is directly-affected by a mechanism, a lack of HTEs could mean that (1) our model of the relationship between $X_{k}$ and mechanism $M$ is wrong; or (2) that no MDV exists because the mechanism is inert or produces a homogeneous effect. Because we cannot rule explanation (1), we cannot affirm explanation (2).
@@
This observation is particularly stark when we consider the statistical properties of HTEs. Low statistical power for interactions reduces our ability to statistically detect HTEs that do exist. In other words, we risk many false-negatives in inferences related to the existence of heterogeneity. Because a lack of HTEs is uninformative about mechanisms, low power suggests that applied researchers often operate in a world in which heterogeneity analysis is unlikely to provide information to support inferences about mechanisms. [^17] Selective reporting of significant results complicates the situation further. In this case, evidence in favor of treatment-effect heterogeneity is more likely to be a false-positive, which increases the the probability that researchers infer that a mechanism is active when it is not.
6.3 Guidance for research design
@@
Our framework posits several recommendations for the design of causal research that seeks to test mechanisms quantitatively using HTE. Our suggestions are premised on improvements in measurement. In terms of covariates, we are primarily concerned with which covariates are measured and the number of candidate MDVs (per mechanism) among those covariates. Covariates are only useful for ascertaining mechanisms when (1) they are plausibly MDVs for a single mechanism; and (2) they do not moderate direct effects. This observation suggests that special care must be taken when positing candidate MDVs. When pre-treatment covariates are (largely) collected in baseline data collection, there is a need to posit MDVs and defend exclusion assumptions ex-ante. Such considerations require more theory and justification than are typically conveyed in the specification of moderation analyses in pre-analysis plans.
@@
When considering directly-affected outcomes, it is very useful to have multiple candidate MDVs for a given mechanism. To see why, consider the case in which we have two candidate MDVs, $X_{1}$ and $X_{2}$ for mechanism $M$, the exclusion assumption holds for both candidate MDVs, and the outcome is directly-affected. Suppose that there do not exist HTEs in $X_{1}$ but there do exist HTEs in $X_{2}$. If we only measured HTEs with respect to $X_{k}$, following Proposition 2, we would not be able to ascertain whether the problem is with the theory ($X_{1}\notin\textbf{X}^{MDV}$) or whether there simply exist no MDV for mechanism $M$. If there exist HTEs in $X_{2}$, we can eliminate the possibility that there do not exist MDV for mechanism $M$. This would suggest that the theory with respect to $X_{k}$ is misspecified. This is useful insofar as it allows us to make an inference that mechanism $M$ is active. Note, however, that in order to leverage multiple candidate MDVs, the exclusion assumption must hold for each candidate MDV, which can be quite demanding.
Our distinction between directly- and indirectly-affected outcomes yields two further recommendations for research design. First, if a goal of a research design is to distinguish causal mechanisms, directly-affected outcomes should be prioritized in HTE analysis. This requires researchers to make clear which outcomes are directly-affected and emphasizes the value of measuring these outcomes. In empirical studies analogous to our motivating example, researchers would typically measure (self-reported) vote choice in a survey to measure the effects of a shock on incumbent support. But if they were running a survey, they could, in principle, elicit willingness-to-pay for the incumbent to try to directly measure voter utility. Our results suggest that the latter would be a worthwhile—if non-standard—investment because HTEs can provide some information about mechanisms with this latter outcome (but not the former).
Second, these results merit broader consideration of latent variable measurement models (Fariss, Kenwick, and Reuning, 2020) in the case when a theorized directly-affected outcome is latent. It is rare for researchers to explicitly measure treatment effects on estimates of a latent variable (e.g., attitudes or preferences). However, popular methods for indexing multiple outcome measures, including $Z$-score indices (i.e., Kling, Liebman, and Katz, 2007), do this implicitly. More explicit consideration of which latent variables are directly affected by (a) mechanism(s) and how may improve the use of HTEs for mechanism detection. Such considerations can also provide information about which latent variable models are most appropriate.
6.4 Strengthening assumptions to use HTE as mechanism tests for indirectly-affected outcomes
Our results suggest that HTEs or lack thereof are uninformative about mechanism activation when measured outcomes are indirectly affected by treatment. It is worthwhile to consider whether we can make progress in this common case by imposing stronger assumptions. To this end, we consider three alternatives. First, we consider an empirical assumption that is widely utilized in partial identification results: monotonicity (Manski, 1997). In our context, monotonicity holds that for all $x^{\prime}>x\in X_{k}$, $CATE(x^{\prime})\geq(\leq)CATE(x)$. Second, we consider the invocation of theoretical assumptions that provide a mapping between an unobserved directly-affected outcome and an observed indirectly-affected outcome. Finally, we ask whether moving from the existence to the magnitude of HTE can generate more information about mechanisms.
Monotonicity: First, in Appendix G.1, we show that the assumption of montonicity is not, in general, sufficient to provide information about mechanism activation through analysis of HTEs. This occurs because an assumption of montonicity fails to provide enough information about the data generating process that generates directly- and indirectly-affected outcomes. We show this in the context of the following oft-used data generating process for a binary outcome, $h(Y)\in\{0,1\}$:
$Y$ $=g(Z,X_{k})+\varepsilon_{i}$
$h(Y)$ =$=\begin{cases}0&\text{ if }Y\leq c\\ 1&\text{ if }Y>c,\end{cases}$
for some constant $c\in(-\infty,\infty)$ and random variable $\varepsilon$ distributed according to the density function $f_{\varepsilon}(\cdot)$. Under the assumption that $g$ is continuous and differentiable, this data-generating process holds that $X_{k}$ is a MDV if $\frac{\partial^{2}g}{\partial Z\partial X_{k}}\neq 0$ for some $x\in X_{k}$. An assumption of monotonicity strengthens this condition by further implying that $\frac{\partial^{2}g}{\partial Z\partial X_{k}}$ is weakly positive or negative for all $x\in X_{k}$. Our goal is to learn whether $X_{k}$ is a MDV by assessing whether $\frac{\partial^{2}g}{\partial Z\partial X_{k}}$ is non-zero by estimating HTEs.
@@ 6.4 Strengthening assumptions to use HTE as mechanism tests for indirectly-affected outcomes
In Proposition 10, we show that montonicity alone is not sufficient to ensure that HTEs take different signs when $X_{k}$ is not a MDV. Moreover, it does not ensure that $\frac{\partial^{2}Y}{\partial Z\partial X_{k}}>(<)0$ when $X_{k}$ is a MDV and montonicity is satisfied. These results show that we need additional assumptions on the distribution $f_{\varepsilon}$ and/or the functional form of $g(Z,X_{k})$ for monotonicity to provide sufficient information to distinguish mechanisms.
Modelling directly-affected outcomes: Second, and in contrast, Appendix G.2 shows how widely-used random utility models do permit inferences about mechanism activation when observed outcomes are indirectly-affected. These models provide a functional mapping between an individual’s utility (the directly-affected outcome) and their choice between relevant alternatives (the indirectly-affected outcome). For example, in our motivating example, a random utility model aims to recover a measure of utility by decomposing a given decision (i.e., vote choice) into an observed systematic component and a random component. By specifying the systematic component as a function of individual- or choice-specific covariates and assuming the distribution of the random component, researchers can estimate the systematic component of utility. Using these estimates of our directly-affected outcome (utility), researchers can then use HTE to assess mechanism activation under Proposition 1. The invocation of a random-utility model is not free: it makes strong assumptions about the mapping between utility and choice outcomes. Researchers may not be well-positioned to invoke or assess these assumptions. However, these assumptions about the mapping between directly- and indirectly-affected outcomes do buy additional information which may permit learning about mechanism activation from HTE.
@@
Learning from the Magnitude of HTEs: To this point, following common practice, we have focused on the existence of HTEs, rather than the magnitude of HTEs. By magnitude of HTEs, we are referring to the the size of the difference in CATEs. In principle, this magnitude may provide additional information. For example, suppose that theory suggested that a given moderator $X_{k}$ may be a MDV for two distinct mechanisms. If this were the case, the exclusion assumptions (Assumptions 1 and 3) would not be satisfied. However, we may have reasons to believe that the HTEs coming from one mechanism are large (in magnitude) while those coming from the other mechanism are small (in magnitude). In Appendix G.3, we propose a simple Bayesian model that allows for inference about the activation of a specified mechanism from the magnitude of estimated effects. This model requires specification of priors over: (1) the probability of activation of a given mechanism and (2) the magnitude of HTEs under each candidate mechanism. Specification of these priors constitutes the invocation of additional assumptions.
@@
We note that most theories in the social sciences admit directional—rather than point—predictions. By moving from the existence of HTE to their magnitude, the model that we propose requires researchers to specify priors about the size of HTEs, not simply their existence or direction. While these priors allow us to glean some information about mechanisms when the exclusion assumption(s) are violated, more work is needed to guide researchers in specifying such priors from theories that are largely directional.
@@ 7 Conclusion
Social scientists routinely estimate HTEs with the goal of understanding which mechanisms generate treatment effects. By providing the first theoretical analysis of the relationship between HTEs and mechanisms, we show that detecting mechanisms with HTEs is far less straightforward than implied by current practice. Specifically, any link between a covariate (moderator) and a mechanism requires an exclusion assumption, so that covariate does not moderate other effects of treatment on an outcome. Even when this assumption holds, we can only use HTEs to affirm the activation of a mechanism when (1) HTEs exist and (2) the outcome is directly affected by a mechanism. Outside this case, HTEs do not provide sufficient information to show that a mechanism is active or inactive. In this sense, HTEs analysis should not be used to “rule out” activation of a mechanism.
At present, mechanism detection is the modal use of HTEs in political science (see Table 1) and the modal method for mechanism detection (Blackwell, Ma, and Opacic, 2024). However, mechanism detection is not the only use of HTE. Our results speak to contexts where mechanistic analysis is a goal. HTEs are also increasingly used for extrapolation of treatment effects to different populations/settings (Egami and Hartman, 2022; Devaux and Egami, 2022) and the targeting of treatments (Athey, Tibshirani, and Wager, 2019; Kitagawa and Tetenov, 2018). Our work does not directly speak to these uses of HTEs, because these methods do not seek to attribute observed effects to mechanisms (Slough and Tyson, 2024a).
Our analysis raises a number of issues and opportunities for future research to build upon. In particular, we emphasize the need to distinguish between the level at which mechanisms operate (e.g., on utility) and the outcomes we observe. This distinction has underappreciated implications for multiple quantitative methods to detect mechanism activation, including mediation analysis, analysis of treatment effects on intermediate outcomes, and efforts to link the sign of treatment effect to a (set of) mechanism(s). One potentially fruitful avenue for continued use of HTEs for mechanism detection would be to move from the presence of HTEs to their magnitude, as we outline in Appendix G.3. Adoption of this approach will rely on the the development of closer links between theoretical mechanisms and reduced-form treatment effects than is current practice.

### v4

@@ 4.2 Improving the interpretation of HTEs as mechanism tests
By summarizing our results, Table 2 provides a guide for interpreting the presence or absence of HTEs as tests of a mechanism. Importantly, they point to an asymmetry in what we can infer about mechanisms from the presence versus absence of HTEs. Specifically, when exclusion assumptions hold, the presence of HTEs provides evidence that a mechanism is active. In contrast, the absence of HTEs does not rule out a mechanism or show that it is inert without further assumptions.
@@
Our results are theoretical or identification results, meaning that they can be interpreted as what would obtain if we had an infinite sample. But, in practice, empiricists operate in a world with finite—and often relatively small—samples. This introduces statistical problems as well. In the context of HTEs, known limits to the statistical power for interaction effects (or terms) reduces our ability to detect HTEs that do exist. In other words, we risk many false-negatives in inferences related to the existence of heterogeneity. Because a lack of HTEs provides less informative about mechanism activation, low power suggests that applied researchers often operate in a world in which heterogeneity analysis is unlikely to provide information to support inferences about mechanisms. [^19] Selective reporting of significant results complicates the situation further. In this case, evidence in favor of treatment-effect heterogeneity is more likely to be a false-positive, which increases the the probability that researchers infer that a mechanism is active when it is not.
In the context of the studies that we examine, authors are admirably transparent about these limitations. For example, de Figueiredo et al. (2023: p. 735) (Panel (d)) write “given the small samples, however, the difference between the two [CATE] estimates (the interaction) is not statistically significant. Still, the difference in magnitudes certainly suggests that [candidate’s] voters are more sensitive to corruption-related information than supporters of other candidates.” In sum, the combination of the asymmetry in the inferences about mechanisms with versus without HTEs combined with the limited ability to detect heterogeneity should give authors caution in the interpretation of HTEs as tests of mechanisms.
4.3 Guidance for prospective research design
Our framework posits several recommendations for the design of causal research that seeks to test mechanisms quantitatively using HTE. Our suggestions are premised on improvements in measurement.
@@ 4.3 Guidance for prospective research design
Measure more candidate MDVs: In terms of covariates, we are primarily concerned with which covariates are measured and the number of candidate MDVs (per mechanism) among those covariates. Following the exclusion assumptions, covariates are only useful for ascertaining mechanisms when (1) they are plausibly MDVs for a single mechanism; and (2) they do not moderate direct effects. This observation suggests that special care must be taken when positing candidate MDVs. When pre-treatment covariates are (largely) collected in baseline data collection, there is a need to posit MDVs and defend exclusion assumptions ex-ante. Such considerations require more theory and justification than are typically conveyed in the specification of moderation or HTE analyses in pre-analysis plans.
@@
Further, it is very useful to have multiple candidate MDVs for a given mechanism. To see why, consider the case in which we have two candidate MDVs, $X_{1}$ and $X_{2}$ for mechanism $M_{1}$ and the exclusion assumptions hold for both candidate MDVs. Suppose that there do not exist HTEs in $X_{1}$ but there do exist HTEs in $X_{2}$. If we only measured HTEs with respect to $X_{1}$, following Proposition 3, we would not be able to ascertain whether the problem is with the theory ($X_{1}\notin\textbf{X}^{MDV}$) or whether there simply exist no MDV for mechanism $M_{1}$. If there exist HTEs in $X_{2}$, we can eliminate the possibility that there do not exist MDV for mechanism $M_{1}$. This would suggest that the theory with respect to $X_{1}$ is misspecified. This is useful insofar as it allows us to make an inference that mechanism $M_{1}$ is active. Note, however, that in order to leverage multiple candidate MDVs, the exclusion assumption must hold for each candidate MDV, which can be quite demanding.
The simplified presentation of the theories in Figure 3 are unlikely to reflect the full set of candidate MDVs for a given mechanism. However, we can explore this logic with respect to the theory in Panel (c) where there are two candidate MDVs. Specifically, in this study, corruption aversion and partisanship are posited to be MDVs for the distaste for corruption mechanism. Anduiza et al. (2013) measure only the partisanship indicator and find that respondent distaste for a (hypothetical) incumbent’s corruption varies in partisanship. Suppose instead that the authors they had also measured corruption aversion—and detected HTEs in that variable—while failing to detect HTEs in partisanship. Such a finding would suggest that the distaste mechanism is active (via the HTEs in corruption aversion). It would additionally indicate that partisanship may not be a MDV for the distaste mechanism, perhaps casting doubt on the role of motivated reasoning in shaping a voters’ distaste.
Prioritize specific outcomes for mechanism tests: Our focus on non-linear transformations of measured outcomes yields a further recommendations for research design. If a goal of a research design is to destinguish between mechanisms or detect a posited mechanism, outcomes for which Assumptions 1 and 2 are plausible should be prioritized in HTE analysis. For example, if researchers had survey measures of voter utility from the incumbent and vote choice (from survey responses of administrative vote returns at the precinct level), any inference about the mechanism should be made on the basis of the utility measure. The logic behind this choice helps to convey the structure of the theorized mechanism and its relationship to the outcome measures. Moreover, ex-ante specification of the outcomes for which these assumptions are likely to hold can guide choices about which outcome variables to invest in measuring. This can also guide pre-analysis plans.
To be clear, our recommendation is not to avoid the estimation of HTEs for non-linearly transformed outcomes entirely. Indeed, in the study of elections, we typically care about vote choice more than voter utility, since this vote choice determines who wins office. But we should be clear about why HTEs on vote choice matter when they are not informative about mechanisms. They could be used to understand how to better target future corruption information interventions across the electorate (e.g., Athey and Wager, 2021; Kitagawa and Tetenov, 2018), to extrapolate effects to other electorates under a model (e.g., Egami and Hartman, 2022), or to describe about the distributional impacts of the treatment. These are all valid—and perhaps underutilized—uses of HTEs which do not depend on the relationship between mechanisms and measured outcomes. We simply advise researchers to more carefully articulate the goals of their analysis.
4.4 Imposing assumptions about to use HTEs as mechanism tests
Our results suggest that the presence of HTEs is not informative of mechanism activation when the exclusion assumptions do not hold. Model-based approaches may be useful when researchers only have access to transformed outcomes or have reasons to doubt the validity of the exclusion assumptions. For example, for the field experimental and observational studies in Figure 3 (Panels (a), (b), and the field experiment in (d)), authors have vote share outcomes or self-reported vote choice, but recall that this is a transformed outcome. Furthermore, in Panel (b), it may also be reasonable to believe that corruption aversion conditions both the distaste for corruption and voter coordination mechanisms. Can we make progress in these settings by imposing stronger assumptions about the mapping from voter preferences (utility) to vote choice?
We consider the possibility of invoking different models or sets of assumptions may aid in using HTEs to learn about mechanisms. Table 3 summarizes three problems identified by our analysis. Each problem is paired with a statistical model or set of assumptions that we examine as a solution. The right column summarizes the conclusions of our analyses, which are detailed in greater depth in H. In sum, two of the three models/sets of added assumptions are strong enough (in isolation) to allow HTE to provide information about mechanism activation, where they do not in the absence of these additional assumptions.
Problem Model/assumptions Conclusions
1 Exclusion assumptions fail because utility is transformed into a discrete choice. Random utlity model specifies a systematic and random component of utility that generates observed outcomes. In settings in which mechanisms operate on utility but we observe choice, random-utility models can recover a model-based estimate of expected utility that can be used for analysis of mechanism activation through HTEs.
2 Exclusion assumptions fail because outcome is transformed. Assumption about monotonicity of CATEs in a candidate MDV $X_{k}$. Monotonicity is not strong enough to provide information about mechanism activation without ancillary functional form assumptions.
3 Exclusion assumptions fail because covariate may moderate multiple mechanisms. Bayesian model of CATE magnitude under different mechanism activation profiles. Plausible when we are willing to register priors about relative CATE magnitudes under different mechanisms.
Table 3: Summary of model- or assumption-based alternatives to exclusion assumptions.
Modelling non-transformed outcomes/random utility models: In the context of transformations from utility to choice—like the transformation of voter utility from the incumbent into a vote for the incumbent—there exist widely-used random utility models that seek to recover preferences (utility) from choices. These models provide a functional mapping between an individual’s utility (the outcome upon which the mechanism operates) and their choice by decomposing vote choice into an observed systematic and an unobserved random component. In our motivating model, the information treatment, the corruption aversion covariate, and a partisan alignment covariate are systematic and could, in principle, be observed. In contrast, the valence shock is random. In Panel (b) of Figure 3, corruption information (the treatment), corruption aversion, and the network structure of a precinct (or a sufficient statistic thereof) could be observed, but partisan alignment is random. By specifying the systematic component as a function of individual- or choice-specific covariates and assuming a distribution of the random component(s), researchers may be able to estimate the systematic component of utility.
Appendix H.2 analyzes the use of random-utility models in the context of HTE analysis. These models have two principal merits in the present context. First, the parameterization of the systematic component allows researchers to assess whether the exclusion assumptions are plausible under a given theoretical model. If these assumptions are plausible, the second benefit of a random utility model emerges. Specifically, it permits researchers to estimate (or “back out”) an outcome for which HTEs can provide information about mechanism activation, specifically $\mathbb{E}[Y]$. Importantly, the invocation of a random-utility model is not free: it makes strong parametric assumptions about utility and its relationship to choice outcomes. Researchers may not be well-positioned to invoke or assess these assumptions. However, these assumptions allow for more formal examination of exclusion assumptions and may yield information which may permit learning about mechanism activation from HTEs.
[…]
Imposing monotonicity: To this point, we analyze the current practice of examining the presence of HTEs as a test of mechanisms. However, examining the magnitudes of estimated CATEs may offer more information. We first consider whether the invocation of monotonicity of CATEs—a common empirical assumption (Manski, 1997)—can provide evidence about mechanisms. In our context, monotonicity holds that for all $x^{\prime}>x\in X_{k}$, $CATE(x^{\prime})\geq(\leq)CATE(x)$. In the case of our model (and the applications in Panels a and b), corruption information should have a stronger (negative) effect on voter utility among voters with stronger corruption aversion (larger $\lambda$). Indeed, Remark 1 shows that for $\lambda>\lambda^{\prime}$, $|CATE(y_{1}|X_{1}=\lambda)|>|CATE(y_{1}|X_{1}=\lambda^{\prime})|$ for the voter utility outcome. Is it possible that this monotonicity in $\lambda$ is maintained for the vote choice outcome? If this were the case, we could simply compare the magnitude of effects at different levels of the MDV ($\lambda)$ for some suggestive evidence about mechanism activation. Similarly, could a violation of monotonicity of the form $\text{sign}(CATE(y_{2}|\lambda))\neq\text{sign}(CATE(y_{2}|\lambda^{\prime}))$ provide evidence against mechanism activation?
@@ 4.4 Imposing assumptions about to use HTEs as mechanism tests
Unfortunately, in Appendix H.1, we show that assuming monotonicity is not sufficient to provide information about mechanism activation through analysis of CATE magnitudes. Specifically, in Proposition 12, we show that montonicity alone is not sufficient to ensure that HTEs take different signs when $X_{k}$ is not a MDV. Moreover, it does not ensure that CATEs are monotonic in $X_{k}$ even when $X_{k}$ is a MDV and montonicity is satisfied for the non-transformed outcome. These results show that we would need additional parametric assumptions for monotonicity to provide sufficient information to distinguish mechanism activation.
@@
Learning from the Magnitude of HTEs: The magnitude of CATEs may provide additional information than the presence of HTE. For example, suppose that—in contrast to Panel (b) of Figure 3—theory suggested that the corruption aversion moderator was a candidate MDV for both the distaste and coordination mechanisms. If this were the case, the exclusion assumptions (Assumptions 1 and 2) would not be satisfied. However, we may have reasons to believe that the HTEs coming from one mechanism are large (in magnitude) while those coming from the other mechanism are small (in magnitude). In Appendix H.3, we propose a simple Bayesian model that allows for inference about the activation of a specified mechanism from the magnitude of estimated effects. This model requires specification of priors over: (1) the probability of activation of a given mechanism and (2) the magnitude of HTEs under each candidate mechanism. Specification of these priors constitutes the invocation of additional assumptions.
@@
We note that most theories in the social sciences admit directional—rather than point—predictions. By moving from the existence of HTE to their magnitude in this Bayesian setting, the model that we propose requires researchers to specify priors about the size of HTEs, not simply their existence or direction. While these priors allow us to glean some information about mechanisms when the exclusion assumption(s) are violated, more work is needed to guide researchers in specifying such priors from theories that are largely directional.
@@ 5 Conclusion
Social scientists routinely estimate HTEs with the goal of understanding which mechanisms generate treatment effects. By providing the first theoretical analysis of the relationship between HTEs and mechanisms, we show that detecting mechanisms with HTEs is far less straightforward than implied by current practice. Specifically, any link between a covariate (moderator) and a mechanism requires exclusion assumptions, so that covariate does not moderate the effects of other mechanisms (or the direct effect). Even when these assumptions hold, we can only use HTEs to affirm the activation of a mechanism when (1) HTEs exist and (2) for transformed outcomes that preserve the additive separability of a mechanism’s effect. Outside this case, HTEs do not provide sufficient information to show that a mechanism is active or inactive. In this sense, HTEs analysis should not be used to rule out activation of a mechanism without stronger assumptions than those that we impose.
At present, mechanism detection is the modal use of HTEs in political science (see Table 1) and the modal method for mechanism detection (Blackwell et al., 2024). However, mechanism detection is not the only use of HTE. Our results speak to contexts where mechanistic analysis is a goal. HTEs are also increasingly used for extrapolation of treatment effects to different populations/settings (Egami and Hartman, 2022; Devaux and Egami, 2022) and the targeting of treatments (Athey et al., 2019; Kitagawa and Tetenov, 2018). Our work does not directly speak to these uses of HTEs, because these methods do not seek to attribute observed effects to mechanisms (Slough and Tyson, 2024).
Our analysis raises a number of issues and opportunities for future research to build upon. In particular, we emphasize that choices about which outcomes we measure can complicate efforts to understand the substantive mechanisms at work. For example, even if the exclusion assumptions hold for one outcome, by imposing a common non-linear transformation on that outcome, the exclusion assumptions can be violated. This distinction has underappreciated implications for multiple quantitative methods to detect mechanism activation, including mediation analysis, analysis of treatment effects on intermediate outcomes, and efforts to link the sign of treatment effect to a (set of) mechanism(s). One potentially fruitful avenue for continued use of HTEs for mechanism detection would be to move from the presence of HTEs to their magnitude, as we outline in Appendix H.3. Adoption of this approach will rely on the the development of closer links between theoretical mechanisms and the size of reduced-form treatment effects than is current practice.